# Towards Efficient Training and Evaluation of Robust Models against $l_{0}$ Bounded Adversarial Perturbations

### v1

@@ 5.1 Evaluation of Different Attack Methods
Figure 1: Comparison between sPGD and RS attack under different iterations. (a) CIFAR-10, $k=20$. ResNet18 (std), PORT ($l_{\infty}$ and $l_{2}$) [39], $l_{1}$-APGD ($l_{1}$) [18] and sTRADES ($l_{0}$) are evaluated. (b) ImageNet-100, $k=200$. ResNet34 (std), Fast-EG-$l_{1}$ ($l_{1}$) [19] and sAT ($l_{0}$) are evaluated. The total iteration number ranges from $20$ and $10000$. The results of sPGD and RS attack are shown in solid lines and dotted lines, respectively. For better visualization, the x-axis is in the log scale.
5.2 Comparison under Different Iteration Numbers
@@
In this subsection, we further compare our method sPGD, which is a white-box attack, with RS attack, the strongest black-box attack in the previous section. Specifically, we compare them under various iteration numbers on CIFAR-10 and ImageNet-100, which have different resolutions. Since CIFAR-100 has the same resolution as CIFAR-10, the results on CIFAR-100 are deferred to Appendix A.2. In addition, we also compare sPGD and RS under different sparsity levels in Appendix A.3.
@@
As illustrated in Figure 1, sTRADES has better performance than other robust models by a large margin in all iterations of both sPGD and RS attacks, which is consistent with the results in Table 12 and 9. For vanilla and other robust models, although the performances of both sPGD and RS attack get improved with more iterations, sPGD outperforms RS attack by a large margin when the iteration number is small (e.g. $<1000$ iterations), which makes it feasible for adversarial training. Similar to other black-box attacks, the performance of RS attack drastically deteriorates when the query budget is limited. In addition, our proposed gradient-based sPGD significantly outperforms RS on ImageNet-100, where the search space is much larger than that on CIFAR-10, i.e., higher image resolution and higher sparsity level $k$. This suggests that our approach is scalable and shows higher efficiency on high-resolution images. Furthermore, although the performance of RS does not converge even when the iteration number reaches $10000$, a larger query budget will make it computationally impractical. Following the setting in Croce et al. [16], we do not consider a larger query budget in our experiments, either. To further showcase the efficiency of our approach, we compare the distributions of the number of iterations needed by sPGD and RS to successfully generate adversarial samples and the runtime of different attacks in Appendix A.4 and A.5, respectively.
@@ 5.3 Transferability of Adversarial Perturbations
To evaluate the transferability of our attack across different models and architectures, we generate adversarial perturbations based on one model and report the attack success rate (ASR) of these perturbations on another model. As shown in Table 3, sPGD exhibits better transferability than the most competitive baseline, Sparse-RS (RS) [16]. This further demonstrates the effectiveness of our method and its potential application in more practical scenarios.
Table 3: Transferability of RS and sPGD between between VGG11 (V) and ResNet18 (R). The results are reported in attack success rate (ASR). The perturbation is generated based on the source model (left), and the ASR is obtained on the target model (right). The evaluated dataset is CIFAR-10 [59], and the sparsity level $k=20$. Note that the models are trained on clean instances, and the step-size of sPGD $\alpha$ and $\beta$ are set to 0.75.
@@
sPGDproj 58.0 43.9 50.8 48.0
sPGDunproj 64.9 40.0 52.7 56.7
5.4 Ablation Studies
We conduct ablation studies in this section. We focus on CIFAR10 and the sparsity level $k=20$. Unless specified, we use the same configurations as in Table 1.
Components of sPGD: We first validate the effectiveness of each component in sPGD. The result is reported in Table 4. We observe that naively decomposing the perturbation $\bm{\delta}$ by $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ and updating them separately can deteriorate the performance. By contrast, the performance significantly improves when we update the mask ${\bm{m}}$ by its continuous alternative $\widetilde{{\bm{m}}}$ and $l_{0}$ ball projection. This indicates that introducing $\widetilde{{\bm{m}}}$ greatly mitigates the challenges in optimizing discrete variables. Moreover, the results in Table 4 indicate the performance can be further improved by the random reinitialization mechanism, which encourages exploration and avoids trapping in a local optimum. In Appendix A.6, we compare the performance when we use different step sizes for the magnitude tensor ${\bm{p}}$ and the sparsity mask ${\bm{m}}$. The results in Table A.6 and A.6 of Appendix A.6 indicate that the performance of our proposed method is quite consistent under different choices of step sizes, which facilities hyper-parameter selection for practitioners.
Table 4: Ablation study of each component in sPGDproj in terms of robust accuracy. The model is trained by Fast-EG-$l_{1}$.
Ablations Robust Acc.
Baseline (PGD0 w/o restart) 49.4
    + Decomposition: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ 58.0 (+8.6)
    + Continuous mask $\widetilde{{\bm{m}}}$ 33.9 (-15.5)
    + Random reinitialization 18.1 (-31.3)
Adversarial training: We conduct preliminary exploration on adversarial training against sparse perturbations, since sPGD can be incorporated into any adversarial training variant. Table 178 study sAT and sTRADES, while sTRADES outperforms sAT in all cases. In addition, the training of sAT is relatively unstable in practice, so we focus on sTRADES for ablation studies in this section. We leave the design of sPGD-adapted adversarial training variants to further improve model robustness as a future work.
Table 5.4 demonstrates the performance when we use different backward functions. The policies include always using the sparse gradient (Proj.), always using the unprojected gradient (Unproj.), alternatively using both backward functions every $5$ epochs (Alter.) and randomly selecting backward functions (Rand.). The results indicate that randomly selecting backward functions has the best performance. In addition, Table 5.4 demonstrates the robust accuracy of models trained by sTRADES with different multi-$k$ strategies. The results indicate that multi-$k$ strategy helps boost the performance. The best robust accuracy is obtained when the adversarial budget for training is $6$ times larger than that for test. Furthermore, we also study the impact of different tolerances during adversarial training in Table A.7 of Appendix A.7. The results show that higher tolerance during adversarial training benefits the robustness of the model.
Policy Proj. Unproj. Alter. Rand.
Acc. 41.5 39.4 51.5 61.7
Table 6: Ablation study on multi-$k$ stretegy during adversarial training. The model is PRN18 trained by sTRADES with random policy. The robust accuracy is obtained through sAA.
$k$ 1×1$\times$1 × 2×2$\times$2 × 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × 10×10$\times$10 ×
Acc. 34.0 39.7 54.5 61.7 60.2 55.5
@@ 6 Conclusion
In this paper, we propose an effective and efficient white-box attack named sPGD to generate sparse perturbations bounded by the $l_{0}$ norm. sPGD obtains the state-of-the-art performance among white-box attacks. Based on this, we combine it with black-box attacks for more comprehensive and reliable $l_{0}$ robustness evaluation. Our proposed sPGD is particularly effective in the realm of limited iteration numbers. Due to its efficiency, we incorporate sPGD into the framework of adversarial training to obtain robust models against sparse perturbations. The models obtained by our training method demonstrate the best robust accuracy.
7 Limitations
During adversarial training, we adopt 20-iter sPGD to generate adversarial examples, so the computational overhead is still high, making it difficult to scale up to super-large datasets. Our subsequent work will focus on developing a faster adversarial training method against $l_{0}$ bounded perturbations without sacrificing its robustness. Furthermore, since our method is evaluated on benchmarks, we do not see it has an obvious negative societal impact.
Acknowledgement
This work is supported by National Natural Science Foundation of China (NSFC Project No. 62306250), CityU APRC Project (Project No. 9610614) and CityU Seed Grant (Project No. 9229130).
Appendix A Additional Experiments
A.1 Results of Different Sparsity Levels and Different Datasets
@@
In this subsection, we present the robust accuracy on CIFAR-10 with the sparsity level $k=10$ and $k=15$, those on CIFAR-100 with $k=10$, those on ImageNet-100 [60] with $k=200$ and those on GTSRB [61] in Table 7, 8 and 9, respectively. It should be noted that since only the training set of GTSRB has annotations, we manually split the original training set into a training set and test set ($1000$ instances). The observations with different sparsity levels and on different datasets are consistent with those in Table 1, which indicates the effectiveness of our method.
@@
Table 7: Robust accuracy of various models on different attacks that generate $l_{0}$ bounded perturbations, where the sparsity level $k=10$. The models are trained on CIFAR-10 [59]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [16]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
@@
Table 8: Robust accuracy of various models on different attacks that generate $l_{0}$ bounded perturbations, where the sparsity level $k=15$. The models are trained on CIFAR-10 [59]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [16]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
@@
Table 9: Robust accuracy of various models on different attacks that generate $l_{0}$ bounded perturbations, where the sparsity level $k=10$. The models are trained on CIFAR-100 [59]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [16]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
@@
A.2 Comparison under Different Iterations on CIFAR-100
Similar to the observation in Section 5.2, Figure 3 also indicates that our method can obtain a higher attack success rate than the strong black-box attack Sparse-RS (RS) when the query budget is limited. Compared with the observation in Figure 1, the efficiency of RS, which is based on heuristic random search, improves when the search space becomes smaller, i.e., smaller image space and lower sparsity level $k$.
A.3 Comparison under Different Sparsity Levels
In this subsection, we compare our method sPGD with RS attack. Specifically, we compare these two attacks under different sparsity levels, i.e., the values of $k$. We can observe from Figure 3 that RS attack shows slightly better performance only on the $l_{\infty}$ model and when $k$ is small. The search space for the perturbed features is relatively small when $k$ is small, which facilitates heuristic black-box search methods like RS attack. As $k$ increases, sPGD outperforms RS attack in all cases until both attacks achieve almost $100\%$ attack success rate.
Figure 2: Comparison between sPGD and RS attack under different iterations on CIFAR-100 with $k=10$. ResNet18 (std), FDA ($l_{\infty}$) [40], $l_{1}$-APGD ($l_{1}$) [18] and sTRADES ($l_{0}$) are evaluated. The total iteration number ranges from $20$ and $10000$. The results of sPGD and RS attack are shown in solid lines and dotted lines, respectively. The x-axis is shown in the log scale.
Figure 3: Comparison between sPGD and RS attack under different sparsity levels. PORT ($l_{\infty}$ and $l_{2}$) [39], $l_{1}$-APGD ($l_{1}$) [18] and sTRADES ($l_{0}$) are evaluated. The $k$ ranges from 00 and $50$. The number of total iterations is set to $10000$. The results of sPGD and RS attack are shown in solid lines and dotted lines, respectively.
A.4 Iteration Numbers Needed to Successfully Generate Adversarial Samples
Figure 4 illustrates the distribution of the iteration numbers needed by sPGD and RS to successfully generate adversarial samples. For $l_{\infty}$, $l_{2}$ and $l_{1}$ robust models, our proposed sPGD consumes distinctly fewer iteration numbers to successfully generate an adversarial sample than RS, the strongest black-box attack in Table 1, while maintaining a high attack success rate. Similar to the observations in Figure 1, the model trained by sTRADES suffers from gradient masking. However, RS still requires a large query budget to successfully generate an adversarial sample. This further demonstrates the efficiency of sPGD, which makes it feasible for adversarial training.
(a) $l_{\infty}$ PORT [39]
(b) $l_{2}$ PORT [39]
(c) $l_{1}$-APGD [18]
(d) sTRADES
Figure 4: Distribution of the iteration numbers needed by sPGD (blue) and RS (orange) to successfully generate adversarial samples. The results are obtained from different models: (a) $l_{\infty}$ PORT [39], (b) $l_{2}$ PORT [39], (c) $l_{1}$-APGD [18] and (d) our sPGD. The average iteration numbers (Avg.) and attack success rate (ASR), i.e., $1-$Robust Acc., are reported in the legend. For better visualization, we clip the minimum iteration number to $10$ and show the x- and y-axis in log scale.
A.5 Runtime of Different Attacks
As shown in Table 10, the proposed sPGD shows the highest efficiency among various attacks. Although sAA consumes more time (approximately 2×2$\times$2 × sPGD + RS), it can provide a reliable evaluation against $l_{0}$ bounded perturbation.
Table 10: Runtime of different attacks on 1000 test instances with batch size 500. The sparsity level $k=20$. The evaluated model is sTRADES. The model is trained on CIFAR-10. The experiments are implemented on NVIDIA Tesla V100.
Attack CS RS SF PGD0 SAIF sPGDproj sPGDunproj sAA
Runtime 604 min 59 min 92 min 750 min 122 min 42 min 45 min 148 min
A.6 More Ablation Study on sPGD
Step Size As shown in Table A.6 and A.6, the robust accuracy does not vary significantly with different step sizes. It indicates the satisfying robustness of our method to different hyperparameter choices. In practice, We set $\alpha$ and $\beta$ to $0.25$ and $0.25\times\sqrt{h\times w}$, respectively. Note that $h$ and $w$ denote the height and width of the image, respectively, which are both $32$ in CIFAR-10.
Tolerance during Attacking As shown in Table A.7, the performance of our method remains virtually unchanged, which showcases that our approach is robust to different choices of tolerance for reinitialization.

### v3

@@ 6.4 Comparison under Different Iteration Numbers and Different Sparsity Levels
Figure 3: Comparison between sPGD and RS attack under different iteration numbers and different sparsity levels. (a) Different iteration number comparison on CIFAR-10, $\epsilon=20$. ResNet18 (std), PORT ($l_{\infty}$ and $l_{2}$) [44], $l_{1}$-APGD ($l_{1}$) [19] and sTRADES ($l_{0}$) are evaluated. (b) Different iteration number comparison on ImageNet-100, $\epsilon=200$. ResNet34 (std), Fast-EG-$l_{1}$ ($l_{1}$) [20] and sAT ($l_{0}$) are evaluated. In (a) and (b), the total iteration number ranges from $20$ and $10000$. For better visualization, the x-axis is in the log scale. (c) Different sparsity comparison on CIFAR-10. The evaluated models are the same as those in (a). The $\epsilon$ ranges from 00 and $50$. The number of total iterations is set to $10000$. Note that the results of sPGD and RS attack are shown in solid lines and dotted lines, respectively.
@@
In this subsection, we further compare our method sPGD, which is a white-box attack, with RS attack, the strongest black-box attack in the previous section. Specifically, we compare them under various iteration numbers on CIFAR-10 and ImageNet-100, which have different resolutions. In addition, we also compare sPGD and RS under different sparsity levels on CIFAR-10.
@@
As illustrated in Figure 3 (a) and (b), sTRADES has better performance than other robust models by a large margin in all iterations of both sPGD and RS attacks, which is consistent with the results in Table 12 and 9. For vanilla and other robust models, although the performances of both sPGD and RS attack get improved with more iterations, sPGD outperforms RS attack by a large margin when the iteration number is small (e.g. $<1000$ iterations), which makes it feasible for adversarial training. Similar to other black-box attacks, the performance of RS attack drastically deteriorates when the query budget is limited. In addition, our proposed gradient-based sPGD significantly outperforms RS on ImageNet-100, where the search space is much larger than that on CIFAR-10, i.e., higher image resolution and higher sparsity level $\epsilon$. This suggests that our approach is scalable and shows higher efficiency on high-resolution images. Furthermore, although the performance of RS does not converge even when the iteration number reaches $10000$, a larger query budget will make it computationally impractical. Following the setting in Croce et al. [17], we do not consider a larger query budget in our experiments, either.
Furthermore, we can observe from Figure 3 (c) that RS attack shows slightly better performance only on the $l_{\infty}$ model and when $\epsilon$ is small. The search space for the perturbed features is relatively small when $\epsilon$ is small, which facilitates heuristic black-box search methods like RS attack. As $\epsilon$ increases, sPGD outperforms RS attack in all cases until both attacks achieve almost $100\%$ attack success rate.
6.5 Efficiency of sPGD
(a) $l_{\infty}$ PORT [44]
(b) $l_{1}$-APGD [19]
(c) sTRADES
Figure 4: Distribution of the iteration numbers needed by sPGD (blue) and RS (orange) to successfully generate adversarial samples. The results are obtained from different models: (a) $l_{\infty}$ PORT [44], (b) $l_{1}$-APGD [19] and (c) our sPGD. The average iteration numbers (Avg.) and attack success rate (ASR), i.e., $1-$Robust Acc., are reported in the legend. For better visualization, we clip the minimum iteration number to $10$ and show the x- and y-axis in log scale.
Table 5: Runtime of different attacks on 1000 test instances with batch size 500. The sparsity level $\epsilon=20$. The evaluated model is sTRADES. The model is trained on CIFAR-10. The experiments are implemented on NVIDIA Tesla V100.
Attack CS RS SF PGD0
Runtime 604 min 59 min 92 min 750 min
Attack SAIF sPGDp sPGDu sAA
Runtime 122 min 42 min 45 min 148 min
To further showcase the efficiency of our approach, we first compare the distributions of the number of iterations needed by sPGD and RS to successfully generate adversarial samples. Figure 4 illustrates the distribution of the iteration numbers needed by sPGD and RS to successfully generate adversarial samples. For $l_{\infty}$ and $l_{1}$ robust models, our proposed sPGD consumes distinctly fewer iteration numbers to successfully generate an adversarial sample than RS, the strongest black-box attack in Table 1, while maintaining a high attack success rate. Similar to the observations in Figure 3, the model trained by sTRADES suffers from gradient masking. However, RS still requires a large query budget to successfully generate an adversarial sample. This further demonstrates the efficiency of sPGD, which makes adversarial training feasible.
Additionally, we compare the runtime of sPGD with other attacks in the same configuration as in Table 1. As shown in Table 5, the proposed sPGD shows the highest efficiency among various attacks. Although sAA consumes more time (approximately 2×2$\times$2 × sPGD + RS), it can provide a reliable evaluation against $l_{0}$ bounded perturbation.
@@ 6.6 Transferability of Adversarial Perturbations
Table 6: Transferability of RS and sPGD between vanilla VGG11 (V) and RN-18 (R) on CIFAR-10 with $\epsilon=20$. Attack success rate (ASR) is reported. The perturbations are generated on the source model (left), and are evaluated on the target model (right). Note that $\alpha$ and $\beta$ of sPGD are set to $0.75$.
@@
sPGDp 58.0 43.9 50.8 48.0
sPGDu 64.9 40.0 52.7 56.7
To evaluate the transferability of our attack across different models and architectures, we generate adversarial perturbations based on one model and report the attack success rate (ASR) on another model. As shown in Table 6, sPGD exhibits better transferability than the most competitive baseline, Sparse-RS (RS) [17]. This further demonstrates the effectiveness of our method and its potential application in more practical scenarios.
6.7 Visualization
(a) $l_{0}$, $\epsilon=200$
(b) $l_{0}$, $\epsilon=200$
(c) row, $\epsilon=1$
(d) row, $\epsilon=1$
(e) star, $25\times 25$, $\epsilon=1$
(f) star, $25\times 25$, $\epsilon=1$
(g) “A”, $60\times 60$, $\epsilon_{\infty}=32/255$
(h) “A”, $60\times 60$, $\epsilon_{\infty}=32/255$
[…]
We show some adversarial examples with different types of sparse perturbations in Figure 5, including $l_{0}$ bounded perturbations, row-wise perturbations, star-like perturbations, and adversarial watermarks of letter “A”. The attack is sPGD, and the model is vanilla ResNet-34 trained on ImageNet-100. More adversarial examples are shown in Appendix D.
@@ 7 Conclusion
In this paper, we propose an effective and efficient white-box attack named sPGD to generate both unstructured and structured perturbations. sPGD obtains the state-of-the-art performance among white-box attacks. Based on this, we combine it with black-box attacks for more comprehensive and reliable evaluation of robustness against sparse perturbations. Our proposed sPGD is particularly effective in the realm of limited iteration numbers. Due to its efficiency, we incorporate sPGD into the framework of adversarial training to obtain robust models against sparse perturbations. The models trained with our method demonstrate the best robust accuracy.
Appendix A Implementation Details
In experiments, we mainly focus on the cases of the sparsity of perturbations $\epsilon=10,15$ and $20$, where $\epsilon=||\sum_{i=1}^{c}\bm{\delta}^{(i)}||_{0}$ or $||{\bm{m}}||_{0}$, $\bm{\delta}^{(i)}\in{\mathbb{R}}^{h\times w}$ is the $i$-th channel of perturbation $\bm{\delta}\in{\mathbb{R}}^{h\times w\times c}$, and ${\bm{m}}\in{\mathbb{R}}^{h\times w\times 1}$ is the sparsity mask in the decomposition of $\bm{\delta}={\bm{p}}\odot{\bm{m}}$, ${\bm{p}}\in{\mathbb{R}}^{h\times w\times c}$ denotes the magnitude of perturbations.
Sparse-PGD (sPGD): Cross-entropy loss is adopted as the loss function of both untargeted and targeted versions of our method. The small constant $\gamma$ to avoid numerical error is set to $2\times 10^{-8}$. The number of iterations $T$ is $10000$ for all datasets to ensure fair comparison among attacks in Table 1. For generating $l_{0}$ bounded perturbations, the step size for magnitude ${\bm{p}}$ is set $\alpha=0.25\times\epsilon_{\infty}$; the step size for continuous mask $\widetilde{{\bm{m}}}$ is set $\beta=0.25\times\sqrt{h\times w}$, where $h$ and $w$ are the height and width of the input image ${\bm{x}}\in{\mathbb{R}}^{h\times w\times c}$, respectively; the tolerance for reinitialization $t$ is set to $3$. For generating structured sparse perturbations, we let $\alpha=0.0125\times\epsilon_{\infty},~{}\beta=0.0125\times\sqrt{h\times w}$ and $t=50$.
Sparse-AutoAttack (sAA): It is a cascade ensemble of five different attacks, i.e., a) untargeted sPGD with unprojected gradient (sPGDu), b) untargeted sPGD with sparse gradient (sPGDp), and c) untargeted Sparse-RS. The hyper-parameters of sPGD are the same as those listed in the last paragraph.
Adversarial Training: sPGD is adopted as the attack during the training phase, the number of iterations is $20$, and the backward function is randomly selected from the two different backward functions for each batch. For sTRADES, we only compute the TRADES loss when training, and generating adversarial examples is based on cross-entropy loss. We use PreactResNet18 [68] with softplus activation [71] for experiments on CIFAR-10 and CIFAR-100 [64], and ResNet34 [67] for experiments on ImageNet100 [65] and GTSRB [66]. We train the model for $100$ epochs on CIFAR-10 and CIFAR-100, for $40$ epochs on ImageNet100, and for $20$ epochs on GTSRB. The training batch size is $128$ on CIFAR-10 and CIFAR-100, and $32$ on ImageNet100 and GTSRB. The optimizer is SGD with a momentum factor of $0.9$ and weight decay factor of $5\times 10^{-4}$. The learning rate is initialized to $0.05$ and is divided by a factor of $10$ at the $\frac{1}{4}$ and the $\frac{3}{4}$ of the total epochs. The tolerance for reinitialization is set to $10$.
Appendix B More Results of Section 6.1
@@ Appendix A Implementation Details
In this subsection, we present the robust accuracy on CIFAR-10 with the sparsity level $\epsilon=10$ and $\epsilon=15$, as well as those on CIFAR-100 with $\epsilon=10$ in Table 7, 8 and 9, respectively. The observations with different sparsity levels and on different datasets are consistent with those in Table 1 and 2, which indicates the effectiveness of our method.
@@
Table 7: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=10$. The models are trained on CIFAR-10 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
@@
Table 8: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=15$. The models are trained on CIFAR-10 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
@@
Table 9: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=10$. The models are trained on CIFAR-100 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
@@
Appendix C Ablation Studies
We conduct ablation studies in this section. We focus on CIFAR10 and the sparsity level $\epsilon=20$. Unless specified, we use the same configurations as in Table 1.